What would be the final temperature if 8.94 x 10^3 joules of heat were added to 454 grams of copper, specific heat 0.386 J/g0C, at 23.00C?
Answer #1

Let the final Temperature be T

Heat added = mass * specific heat capacity of copper * Change in temperature

8.94 * 10^3 = 454 * 0.386 * (T-23.00)

(T-23.00) = 51.01

T = 23.00 + 51.01 = 74.01

Hence the final temperature will be equal to 74.01 Celsius

Note - Post any doubts/queries in comments section. Rate the answer positive if found helpful